OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 20°C and 160°C. The temperature is at its lowest point when t = 0, and the reaction completes one cycle over an eight-hour period. What is a cosine function that would model this reaction?

OpenStudy (anonymous):

A. f(t) = -90 cos pi over 4t + 70 B. f(t) = -70 cos pi over 4t + 90 C. f(t) = 70 cos 8t + 90 D. f(t) = 90 cos 8t + 70

OpenStudy (anonymous):

I'm thinking it's C because I believe that the constant is 70?

OpenStudy (anonymous):

@ParthKohli @wade123

OpenStudy (anonymous):

@zoomiemoo

OpenStudy (anonymous):

Okay it wasn't C

OpenStudy (anonymous):

sorry @twinkle12377 Im not very good as trig I am better with algebra

OpenStudy (anonymous):

I can usually help with calculus as well

OpenStudy (anonymous):

personally I would graph the functions and choose the one the is 20 at 0 and 160 at 8

OpenStudy (anonymous):

ohhh okay thank you so much! :) @zoomiemoo